# The Inadequacy of Similarity-based Privacy Metrics:

### v4

@@ 1 Introduction
The established framework to limit information leakage from trained models is Differential Privacy (DP) [32, 34]; in the context of synthetic data, generative models should be trained while satisfying DP end to end [139, 68, 75]. However, while several synthetic data providers claim their products adhere to regulations like GDPR, HIPAA, or CCPA (see Sec. 2.2), many of them use ad-hoc heuristics to demonstrate privacy empirically rather than DP. Some combine DP with unperturbed heuristics, which breaks the end-to-end DP pipeline and ultimately negates its privacy guarantees, as our analysis confirms. This is worrisome, as these products are typically trained on sensitive data and are already deployed in highly regulated environments, e.g., healthcare [63].
@@
Countermeasures (or Lack Thereof). As mentioned, providing meaningful privacy guarantees would entail training the generative models while satisfying DP. However, we also demonstrate that DP models are still vulnerable to our attacks if used in conjunction with unperturbed heuristics, as the leakage persists through the metrics (see Sec. 8.2).
@@
    We present a simple attack, DifferenceAttack, which achieves perfect performance for membership and attribute inference with just a handful of metrics API calls; see the orange/purple bars in Fig. 1.
@@
    We introduce ReconSyn, a novel reconstruction attack with minimal assumptions, i.e., black-box access to a single trained generative model and the privacy metrics. After applying both attack phases, SampleAttack and SearchAttack, ReconSyn reconstructs 78%–100% of the outliers in all datasets; see the blue bars in Fig. 1.
@@ 2.2 Commercial Synthetic Data Solutions
Company Claimed DP Ad-hoc
Compliance Models Privacy Heuristics
@@
Mostly AI [82] IMS, DCR, NNDR [80]
Hazy2 [56] DCR [58]
@@
Replica3 [103] SF [102]
Statice4 [112] DCR5 [111]
1Acquired by NVIDIA in Mar. 2025.  2Acquired by SAS in Nov. 2024.  3Acquired by Aetion in Jan. 2022.  4Acquired by Anonos in Nov. 2022.  5Used for linkability and inference risk metrics [45].
Table 1: Synthetic data companies and whether they claim to be offering regulatory-compliant synthetic data or DP models, as well as the similarity-based privacy metrics/filters they use.
We have been tracking the main synthetic data providers (Gretel, Tonic, Mostly AI, Hazy, Aindo, DataCebo, Syntegra, YData, Synthesized, Syntho, Replica Analytics, and Statice) and examined publicly available information on their claims of regulatory compliance and their approach to privacy, i.e., whether they support DP training and what heuristics they use. As summarized in Table 1, they mainly use five heuristics, which we describe in Sec. 2.3: three metrics, Identical Match Share (IMS), Distance to Closest Records (DCR), Nearest Neighbor Distance Ratio (NNDR), and two filters, Similarity Filter (SF) and Outlier Filter (OF). N.B. For simplicity, we refer to the three metrics as similarity-based privacy metrics (SBPMs) and use the term (ad-hoc) heuristics to include all metrics and filters.
Almost all companies claim on their websites that their synthetic data products comply with regulations like GDPR, HIPAA, CCPA, etc., even though there are no established standards for mapping privacy to regulatory frameworks in the context of synthetic data. We also note that larger and better-funded companies tend to report adopting DP. This might not be surprising, as integrating DP in production requires specialized technical knowledge. Six of the twelve companies do not support DP but claim to guarantee privacy through one or more of the three SBPMs, while two combine it with the two privacy filters.
@@ 4 Counter-Examples
Next, we discuss three counter-examples whereby there is obvious privacy leakage even if all three tests from SBPMs pass (or any of the privacy filters are applied). We use 2d Gauss dataset (see Appendix B), which we visualize over its two dimensions in Fig. 4. Since all attributes are continuous, we use the Euclidean distance to make the computations more accurate. In Appendix D, we discuss three additional counter-examples, further showcasing the inconsistent nature of the metrics and filters.
@@
CE2. Leaking All Train Outliers. Next, assume that the synthetic data contains all train outliers (with an indistinguishably small perturbation) and the value (0, 0) repeated five times the size of the train data, as displayed in Fig. 4. Again, all privacy tests pass: there are no exact matches, and even though the synthetic outliers are extremely close to the train ones, the large sample size of zeros skews the distances enough to fool both DCR and NNDR. If this synthetic dataset is released, individuals whose data corresponds to the outliers and whose sensitive attributes are leaked would be unconvinced that their privacy is actually preserved [90].
CE3. SF & Reconstruction. We can reconstruct all outliers if Similarity Filter (SF) is applied. We assume access to an oracle possessing knowledge of the generative process. Suppose we use the oracle to sample 100,000 synthetic datasets, apply SF, and select the datasets passing all privacy tests (we plot them in Fig. 4). Since no generative model was trained (i.e., the train data was never exposed to a model), any data directly sampled from the oracle preserves privacy.
@@ 5.1 Adversarial Model
2) Adding/removing records to the synthetic data. The adversary can conditionally generate data or update the synthetic data with desirable patterns (e.g., bias correction, rebalancing, signal boosting) as this does not have an effect on the train data privacy as long as all privacy tests pass (see I2 in Sec. 3). Data augmentation/conditional generation is also a popular use case advertised by synthetic data companies [121, 48, 81, 103].
3) Querying the privacy metrics. The adversary can run the privacy metrics and access the scores if all tests pass. For companies that rely on empirical metrics, privacy is considered a property of the data (see I5 in Sec. 3); thus, querying the metrics on each synthetic dataset and providing access to the scores to their users becomes essential to providing privacy assurances. To be precise, this is explicitly offered by companies [80, 133, 118]; basically, it is their only way to “demonstrate” privacy to users. Prior work on attacking deployed systems also use (large) amounts of queries – e.g., adversaries targeting large language models can query them up to 400,000,000 times to steal parts of the model [16], generate up to 20,000,000,000 tokens to extract portions of the train data [84], etc.
@@ 5.3 Reconstruction Attack
Next, we present ReconSyn, a reconstruction attack aimed at recovering the outliers in the train data. An overview of the attack is shown in Fig. 5. Reconstruction is a more challenging and impactful adversarial task than membership/attribute inference, as it does not assume access to a target record, yet, the adversary can recover entire records from the train data.
@@
Following the second counter-example in Sec. 4, the aggregate scores can be tricked to expose exact distances. Specifically, we can determine the distance between a target synthetic record and the nearest train data counterpart by conditionally generating or augmenting the input synthetic data. One effective way is to submit the target point alongside, e.g., 100 copies of a frequently appearing record whose closest distance we know. This essentially simulates a scenario where the metrics API reveals individual distances to all submitted data (which we do in lines 14, 26, and 33), allowing the detection of matches with 100% confidence.
@@
Efficiency. One could argue that using the Hamming distance limits the adversary’s efficiency, as it does not provide any sense of direction (at all times, any value is either an exact match or not). Nonetheless, our attack achieves strong performance in reconstructing the train outliers (as shown in Sec. 6) and is computationally practical. In all our experimental settings, every phase runs in less than 24 hours on an m4.4xlarge (16 CPUs, 64GB RAM) AWS instance.
@@ 6.1.1 Membership Inference
DifferenceAttack achieves perfect results vs. all generative models and takes under a second (see Table 2) by making just two calls to the privacy metrics, as discussed in Sec. 5.2. Apart from the target record $r_{t}$ (the complete record, including all attributes), the adversary does not have access to any side information.
@@ 6.1.2 Attribute Inference
As discussed in Sec. 5.2, DifferenceAttack can perfectly (AUC score = 1.0) predict the unknown attribute of a partial train record with $k$ calls to the metrics, where $k$ is the number of distinct categories of the unknown attribute. This process takes just a few seconds.
@@ 6.2.2 SampleAttack with Multiple Calls
2d Gauss. For 2d Gauss, we attack the ‘oracle’ (introduced in CE3 in Sec. 4), which possesses knowledge of the generative process. Even though no generative model has been exposed to train data, and the oracle has no memory of the synthetic data it generated, SampleAttack reconstructs 95% of train outliers due to leakage from the privacy metrics.
Adult Small. For Adult Small, we use SampleAttack against all five generative models and report the number of reconstructed outliers in Fig. 6 (top five lines). For PrivBayes, MST, and PATE-GAN, the attack quickly reconstructs about 90% outliers after just ten rounds and eventually reaches 100%. For DPGAN and CTGAN, the attack plateaus at around 85% after 40 rounds, but by round 1,000, it improves and achieves 96% and 99%, respectively. We believe that SampleAttack is extremely successful on this dataset because of its small cardinality ($10^{5}$).
Adult. The models are much less likely to memorize and reproduce individual data points on Adult, which has twice the number of columns and cardinality of $10^{15}$. Indeed, apart from PrivBayes, SampleAttack only recovers 5% of the outliers – see Table 3 and Fig. 6 (bottom five lines).
Census. Even though Census has roughly twice the columns/rows and much higher cardinality, SampleAttack is more successful (excluding PrivBayes on Adult), recovering on average 53% outliers (see middle lines in Fig. 6). Interestingly, attacking CTGAN yields better results than PrivBayes. Also, the recovery rate follows a linear trend (vs. logarithmic for Adult Small).
@@
MNIST. Finally, looking at MNIST, which has even higher dimensionality/cardinality, SampleAttack fails completely and does not reconstruct even a single outlier. In fact, in Fig. 7, we see that CTGAN and PrivBayes generate images with the highest similarity to the real ones but still at a Hamming distance of at least 6 (i.e., number of different pixels). All models, however, create outliers further away from the train data compared to the distances between test and train. This confirms that all privacy tests pass and that test/synthetic datasets do not contain train data copies.
@@ 6.3 SearchAttack
MNIST. Despite MNIST’s high cardinality, SearchAttack reconstructs more than 80% of the train outliers. To reduce the search space, the adversary can be strategic, e.g., excluding some pixels (e.g., the ones on the sides of the image or the ‘frame’) by setting their value to 0 after observing the common pattern and generating a collection of potential outliers. This restricts the adversary’s capability, and they cannot fully reconstruct 21 out of the 488 outliers. Nonetheless, this could be considered a good tradeoff vis-à-vis the number of saved computations, a factor of $480=30\cdot 16$ (30 fixed pixels, 16 combinations per bin) per search. In the last column in Table 3, we report the number of train outliers reconstructed exactly and those within 1 pixel (even though the adversary can easily get an exact train data match by running the attack for one step without any restrictions). A subset of the recovered digits is shown in Fig. 8.
Overall, SearchAttack is very successful, despite SampleAttack’s failure to recover any outliers. Attacking all models, except for CTGAN, results in reconstructing at least 97% of outliers. We believe this high score is due to the generators’ ability to create diverse images not too dissimilar from the outliers (as shown in Fig. 7). Conversely, although CTGAN generates the closest images, that does not recover more outliers. This might be due to mode collapse or its embedding strategy for categorical columns (both DPGAN and PATE-GAN use simple one-hot encoding). Unsurprisingly, out of the restricted 21 outliers, attacking CTGAN leads to recovering only six compared to at least 16 for the other models.
@@ 7 Related Work
Stadler et al. [110] present a systematic evaluation of MIAs and AIAs in the context of tabular synthetic data and show that outliers are vulnerable when model are trained without DP guarantees. Annamalai et al. [6] introduce a new attribute inference attack based on linear reconstruction. Overall, as discussed in Sec. 5.2, unlike state-of-the-art MIAs/AIAs, DifferenceAttack requires no reference data and relies on a single pre-trained model.
@@ 8.1 Summary
Responsible Disclosure. Our main goal is not to attack live systems or access any personal data but to demonstrate the need for formal notions of privacy when rolling out systems in production. We shared our work with the two main synthetic data companies using SBPMs, providing them with 90 days for a response while keeping the paper confidential. To the best of our knowledge, one company stopped offering privacy filters alongside DP model training while the other updated its marketing materials.
@@ 8.2.1 DP Training
We report the resulting privacy-utility tradeoffs in Fig. 9. As in prior work [119], we measure utility through similarity, reporting the average difference between all 1-way marginals and 2-way mutual information scores (normalized in [0,1]) across train and synthetic data. As expected, applying DP reduces utility across the board. The utility of MST drops less than PrivBayes, and similarly, PATE-GAN less than DPGAN, due to the different underlying DP mechanisms, as pointed out in [42].
@@
We measure privacy in terms of the proportion of train outliers reconstructed by ReconSyn. Regardless of the model, privacy budget, or utility, we successfully recover more than 95% of its targets (to help visualize this, we add a dashed vertical line to Fig. 9). Applying DP to higher utility models – i.e., MST and PATE-GAN – does not even defend against SampleAttack, while it yields a big drop in utility for PrivBayes and DPGAN. Nonetheless, SearchAttack also recovers all train outliers against these two models.
Training models with DP guarantees does not help because the leakage comes from the privacy metrics; since they require access to the train data and are deterministic (as discussed in I4 in Sec. 3), they break the end-to-end DP pipeline. Any additional privacy mechanism applied on top of the metrics is unlikely to mitigate the problem, as the analysis of the privacy guarantees of the whole system needs careful consideration [26].

### v5

@@ 1 Introduction
The established framework to limit information leakage from trained models is Differential Privacy (DP) [32, 34]; in the context of synthetic data, generative models should be trained while satisfying DP end to end [139, 68, 75]. However, while several synthetic data providers claim their products adhere to regulations like GDPR, HIPAA, or CCPA (see Section 2.2), many of them use ad-hoc heuristics to demonstrate privacy empirically rather than DP. Some combine DP with unperturbed heuristics, which breaks the end-to-end DP pipeline and ultimately negates its privacy guarantees, as our analysis confirms. This is worrisome, as these products are typically trained on sensitive data and are already deployed in highly regulated environments, e.g., healthcare [63].
@@
Countermeasures (or Lack Thereof). As mentioned, providing meaningful privacy guarantees would entail training the generative models while satisfying DP. However, we also demonstrate that DP models are still vulnerable to our attacks if used in conjunction with unperturbed heuristics, as the leakage persists through the metrics (see Section 8.2).
@@
    We present a simple attack, DifferenceAttack, which achieves perfect performance for membership and attribute inference with just a handful of metrics API calls; see the orange/purple bars in Figure 1.
@@
    We introduce ReconSyn, a novel reconstruction attack with minimal assumptions, i.e., black-box access to a single trained generative model and the privacy metrics.After applying both attack phases, SampleAttack and SearchAttack, ReconSyn reconstructs 78%–100% of the outliers in all datasets; see the blue bars in Figure 1.
@@ 2.2 Commercial Synthetic Data Solutions
Company Claimed DP Ad-hoc Privacy
Compliance Models Heuristics
@@
Mostly AI [82] 2 IMS, DCR, NNDR [80]
Hazy3 [56] DCR [58]
@@
Replica4 [103] SF [102]
Statice5 [112] DCR6 [111]
1Acquired by NVIDIA in Mar 2025. 2DP support, alongside SBPMs, added in Nov 2024. 3Acquired by SAS in Nov 2024. 4Acquired by Aetion in Jan 2022. 5Acquired by Anonos in Nov 2022. 6Used for linkability and inference risk metrics [45].
Table 1: List of representative synthetic data companies, whether they claim to be offering regulatory-compliant synthetic data, and the similarity-based privacy metrics/filters they use.
We have been tracking the main synthetic data providers (Gretel, Tonic, Mostly AI, Hazy, Aindo, DataCebo, Syntegra, YData, Synthesized, Syntho, Replica Analytics, and Statice) and examined publicly available information on their claims of regulatory compliance and their approach to privacy, i.e., whether they support DP training and what heuristics they use. As summarized in Table 1, they mainly use five heuristics, which we describe in Section 2.3: three metrics, Identical Match Share (IMS), Distance to Closest Records (DCR), Nearest Neighbor Distance Ratio (NNDR), and two filters, Similarity Filter (SF) and Outlier Filter (OF). N.B. For simplicity, we refer to the three metrics as similarity-based privacy metrics (SBPMs) and use the term (ad-hoc) heuristics to include all metrics and filters.
Almost all companies claim on their websites that their synthetic data products comply with regulations like GDPR, HIPAA, CCPA, etc., even though there are no established standards for mapping privacy to regulatory frameworks in the context of synthetic data. We also note that larger and better-funded companies tend to report adopting DP. This might not be surprising, as integrating DP in production requires specialized technical knowledge. Five of the twelve companies do not support DP but claim to guarantee privacy through one or more of the three SBPMs, while two combine it with the two privacy filters.
@@ 4 Counter-Examples
Next, we discuss three counter-examples whereby there is obvious privacy leakage even if all three tests from SBPMs pass (or any of the privacy filters are applied). We use 2d Gauss dataset (see Appendix B), which we visualize over its two dimensions in Figure 4. Since all attributes are continuous, we use the Euclidean distance to make the computations more accurate. In Appendix D, we discuss three additional counter-examples, further showcasing the inconsistent nature of the metrics and filters.
@@
CE2. Leaking All Train Outliers. Next, assume that the synthetic data contains all train outliers (with an indistinguishably small perturbation) and the value (0, 0) repeated five times the size of the train data, as displayed in Figure 4. Again, all privacy tests pass: there are no exact matches, and even though the synthetic outliers are extremely close to the train ones, the large sample size of zeros skews the distances enough to fool both DCR and NNDR. If this synthetic dataset is released, individuals whose data corresponds to the outliers and whose sensitive attributes are leaked would be unconvinced that their privacy is actually preserved [90].
CE3. SF & Reconstruction. We can reconstruct all outliers if Similarity Filter (SF) is applied. We assume access to an oracle possessing knowledge of the generative process. Suppose we use the oracle to sample 100,000 synthetic datasets, apply SF, and select the datasets passing all privacy tests (we plot them in Figure 4). Since no generative model was trained (i.e., the train data was never exposed to a model), any data directly sampled from the oracle preserves privacy.
@@ 5.1 Adversarial Model
2) Adding/removing records to the synthetic data. The adversary can conditionally generate data or update the synthetic data with desirable patterns (e.g., bias correction, rebalancing, signal boosting) as this does not have an effect on the train data privacy as long as all privacy tests pass (see I2 in Section 3). Data augmentation/conditional generation is also a popular use case advertised by synthetic data companies [121, 48, 81, 103].
3) Querying the privacy metrics. The adversary can run the privacy metrics and access the scores if all tests pass. For companies that rely on empirical metrics, privacy is considered a property of the data (see I5 in Section 3); thus, querying the metrics on each synthetic dataset and providing access to the scores to their users becomes essential to providing privacy assurances. To be precise, this is explicitly offered by companies [80, 133, 118]; basically, it is their only way to “demonstrate” privacy to users. Prior work on attacking deployed systems also use (large) amounts of queries – e.g., adversaries targeting large language models can query them up to 400,000,000 times to steal parts of the model [16], generate up to 20,000,000,000 tokens to extract portions of the train data [84], etc.
@@ 5.3 Reconstruction Attack
Next, we present ReconSyn, a reconstruction attack aimed at recovering the outliers in the train data. An overview of the attack is shown in Figure 5. Reconstruction is a more challenging and impactful adversarial task than membership/attribute inference, as it does not assume access to a target record, yet, the adversary can recover entire records from the train data.
@@
Following the second counter-example in Section 4, the aggregate scores can be tricked to expose exact distances. Specifically, we can determine the distance between a target synthetic record and the nearest train data counterpart by conditionally generating or augmenting the input synthetic data. One effective way is to submit the target point alongside, e.g., 100 copies of a frequently appearing record whose closest distance we know. This essentially simulates a scenario where the metrics API reveals individual distances to all submitted data (which we do in lines 14, 26, and 33), allowing the detection of matches with 100% confidence.
@@
Efficiency. One could argue that using the Hamming distance limits the adversary’s efficiency, as it does not provide any sense of direction (at all times, any value is either an exact match or not). Nonetheless, our attack achieves strong performance in reconstructing the train outliers (as shown in Section 6) and is computationally practical. In all our experimental settings, every phase runs in less than 24 hours on an m4.4xlarge (16 CPUs, 64GB RAM) AWS instance.
@@ 6.1.1 Membership Inference
DifferenceAttack achieves perfect results vs. all generative models and takes under a second (see Table 2) by making just two calls to the privacy metrics, as discussed in Section 5.2. Apart from the target record $r_{t}$ (the complete record, including all attributes), the adversary does not have access to any side information.
@@ 6.1.2 Attribute Inference
As discussed in Section 5.2, DifferenceAttack can perfectly (AUC score = 1.0) predict the unknown attribute of a partial train record with $k$ calls to the metrics, where $k$ is the number of distinct categories of the unknown attribute. This process takes just a few seconds.
@@ 6.2.2 SampleAttack with Multiple Calls
2d Gauss. For 2d Gauss, we attack the ‘oracle’ (introduced in CE3 in Section 4), which possesses knowledge of the generative process. Even though no generative model has been exposed to train data, and the oracle has no memory of the synthetic data it generated, SampleAttack reconstructs 95% of train outliers due to leakage from the privacy metrics.
Adult Small. For Adult Small, we use SampleAttack against all five generative models and report the number of reconstructed outliers in Figure 6 (top five lines). For PrivBayes, MST, and PATE-GAN, the attack quickly reconstructs about 90% outliers after just ten rounds and eventually reaches 100%. For DPGAN and CTGAN, the attack plateaus at around 85% after 40 rounds, but by round 1,000, it improves and achieves 96% and 99%, respectively. We believe that SampleAttack is extremely successful on this dataset because of its small cardinality ($10^{5}$).
Adult. The models are much less likely to memorize and reproduce individual data points on Adult, which has twice the number of columns and cardinality of $10^{15}$. Indeed, apart from PrivBayes, SampleAttack only recovers 5% of the outliers – see Table 3 and Figure 6 (bottom five lines).
Census. Even though Census has roughly twice the columns/rows and much higher cardinality, SampleAttack is more successful (excluding PrivBayes on Adult), recovering on average 53% outliers (see middle lines in Figure 6). Interestingly, attacking CTGAN yields better results than PrivBayes. Also, the recovery rate follows a linear trend (vs. logarithmic for Adult Small).
@@
MNIST. Finally, looking at MNIST, which has even higher dimensionality/cardinality, SampleAttack fails completely and does not reconstruct even a single outlier. In fact, in Figure 7, we see that CTGAN and PrivBayes generate images with the highest similarity to the real ones but still at a Hamming distance of at least 6 (i.e., number of different pixels). All models, however, create outliers further away from the train data compared to the distances between test and train. This confirms that all privacy tests pass and that test/synthetic datasets do not contain train data copies.
@@ 6.3 SearchAttack
MNIST. Despite MNIST’s high cardinality, SearchAttack reconstructs more than 80% of the train outliers. To reduce the search space, the adversary can be strategic, e.g., excluding some pixels (e.g., the ones on the sides of the image or the ‘frame’) by setting their value to 0 after observing the common pattern and generating a collection of potential outliers. This restricts the adversary’s capability, and they cannot fully reconstruct 21 out of the 488 outliers. Nonetheless, this could be considered a good tradeoff vis-à-vis the number of saved computations, a factor of $480=30\cdot 16$ (30 fixed pixels, 16 combinations per bin) per search. In the last column in Table 3, we report the number of train outliers reconstructed exactly and those within 1 pixel (even though the adversary can easily get an exact train data match by running the attack for one step without any restrictions). A subset of the recovered digits is shown in Figure 8.
Overall, SearchAttack is very successful, despite SampleAttack’s failure to recover any outliers. Attacking all models, except for CTGAN, results in reconstructing at least 97% of outliers. We believe this high score is due to the generators’ ability to create diverse images not too dissimilar from the outliers (as shown in Figure 7). Conversely, although CTGAN generates the closest images, that does not recover more outliers. This might be due to mode collapse or its embedding strategy for categorical columns (both DPGAN and PATE-GAN use simple one-hot encoding). Unsurprisingly, out of the restricted 21 outliers, attacking CTGAN leads to recovering only six compared to at least 16 for the other models.
@@ 7 Related Work
Stadler et al. [110] present a systematic evaluation of MIAs and AIAs in the context of tabular synthetic data and show that outliers are vulnerable when model are trained without DP guarantees. Annamalai et al. [6] introduce a new attribute inference attack based on linear reconstruction. Overall, as discussed in Section 5.2, unlike state-of-the-art MIAs/AIAs, DifferenceAttack requires no reference data and relies on a single pre-trained model.
@@ 8.1 Summary
Responsible Disclosure. Our main goal is not to attack live systems or access any personal data but to demonstrate the need for formal notions of privacy when rolling out systems in production. We shared our work with the two main synthetic data companies using SBPMs, providing them with 90 days for a response while keeping the paper confidential. Since then, one company stopped offering privacy filters alongside DP model training while the other started offering DP model support alongside SBPMs and updated its marketing materials.
@@ 8.2.1 DP Training
We report the resulting privacy-utility tradeoffs in Figure 9. As in prior work [119], we measure utility through similarity, reporting the average difference between all 1-way marginals and 2-way mutual information scores (normalized in [0,1]) across train and synthetic data. As expected, applying DP reduces utility across the board. The utility of MST drops less than PrivBayes, and similarly, PATE-GAN less than DPGAN, due to the different underlying DP mechanisms, as pointed out in [42].
@@
We measure privacy in terms of the proportion of train outliers reconstructed by ReconSyn. Regardless of the model, privacy budget, or utility, we successfully recover more than 95% of its targets (to help visualize this, we add a dashed vertical line to Figure 9). Applying DP to higher utility models – i.e., MST and PATE-GAN – does not even defend against SampleAttack, while it yields a big drop in utility for PrivBayes and DPGAN. Nonetheless, SearchAttack also recovers all train outliers against these two models.
Training models with DP guarantees does not help because the leakage comes from the privacy metrics; since they require access to the train data and are deterministic (as discussed in I4 in Section 3), they break the end-to-end DP pipeline. Any additional privacy mechanism applied on top of the metrics is unlikely to mitigate the problem, as the analysis of the privacy guarantees of the whole system needs careful consideration [26].